OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12

jagr2713 (jagr2713):

i think you use the slope formula

OpenStudy (anonymous):

can you walk me through it @jagr2713

OpenStudy (anonymous):

ANSWER CHOICES A. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9 B. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 D. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3

OpenStudy (anonymous):

the slope formula is y=mx+b you need to put it in that form

OpenStudy (anonymous):

add 11x to both sides now itll be 9y+11x-12

OpenStudy (anonymous):

divide all by 9 now the equation is y=11/9-12/9 can 12/9 be lowered?

OpenStudy (anonymous):

uhM.....I believe so

OpenStudy (anonymous):

what can it be reduced to?

OpenStudy (anonymous):

4/3?

OpenStudy (anonymous):

positive or negative?

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

it would be negative because the 12/9 was negative

OpenStudy (anonymous):

y=mx+b do you know what the slope is?

OpenStudy (igreen):

y = mx + b Where m = slope And b = y-intercept

OpenStudy (igreen):

So if our equation is in the form of y = mx + b. What is the slope and what is the y-intercept? y = 11/9x - 4/3

OpenStudy (anonymous):

11/9 and 4/3?

OpenStudy (anonymous):

is 4/3 positive or negative?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

yes so the slope is 11/9 and y intercept is -4/3

OpenStudy (anonymous):

awesm

OpenStudy (anonymous):

AWESOME THANKS GUYS